a sample of hydrogen gas is mixed with water vapor. the mixture has a total pressure of 755 torr and the water vapor has a partial pressure of 24 torr. what amount (in moles) of hydrogen gas is contained in 1.55 L of this mixture at 298 K?
Answer #1

Ptotal= PPH2 + PPH2O---> PPH2= 755-24=731torr

PV=nRT

760 torr ----1 atm

731 torr -----x= 0.962 atm

n= 0.962 x 1.55/0.082x298= 0.061 mol H2